Diffusion Morning Report 06/14/01

ted berg tberg at snf.stanford.edu
Mon Jun 14 06:33:28 PDT 2004


Good Morning All.
    Looking Pretty good this morning.  Heated Gate valve was installed 
and tested FRiday.  Nitide still under observation.  TEOS 2 is testing.
Everything else looking OK for now.  ted

Diffusion 	Date: 	6/14/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing-recipe issues 	 
Tylan 8-SiGe 	Up 	  	7.29/30
Tylan 9-Poly 	Up 	  	5.19/30
Tylan 10-Nitride 	yellow 	Particles? 	19.35/30
Tylan 11-BPSG-LTO 	Up 	  	13.24/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	13.25/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 440 PSI


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040614/cb9a4a35/attachment.html>


More information about the diffusion mailing list